OpenStudy (anonymous):

FAN AND MEDAL !! Find the derivative of f(x)=-7x^2+5x and f(x)= 4x-8

OpenStudy (anonymous):

see do you remember the identity that: \[\frac{ d }{ dx } x^n = nx^{n-1}\] by applying the same to above will leas us to the answer: \[\frac{ d }{ dx} 7x^2 + 5x = 5 - 14x\] and the same will apply to other EQN. too: \[\frac{ d }{ dx }4x - 8 = 4\]

OpenStudy (anonymous):

@Reggie97g